# InvFussion: Bridging Supervised and Zero-shot Diffusion for Inverse Problems

### v1

@@ Abstract
Diffusion Models have demonstrated remarkable capabilities in handling inverse problems, offering high-quality posterior-sampling-based solutions. Despite significant advances, a fundamental trade-off persists, regarding the way the conditioned synthesis is employed: Training-based methods achieve high quality results, while zero-shot approaches trade this with flexibility. This work introduces a framework that combines the best of both worlds – the strong performance of supervised approaches and the flexibility of zero-shot methods. This is achieved through a novel architectural design that seamlessly integrates the degradation operator directly into the denoiser. In each block, our proposed architecture applies the degradation operator on the network activations and conditions the output using the attention mechanism, enabling adaptation to diverse degradation scenarios while maintaining high performance. Our work demonstrates the versatility of the proposed architecture, operating as a general MMSE estimator, a posterior sampler, or a Neural Posterior Principal Component estimator. This flexibility enables a wide range of downstream tasks, highlighting the broad applicability of our framework. The proposed modification of the denoiser network offers a versatile, accurate, and computationally efficient solution, demonstrating the advantages of dedicated network architectures for complex inverse problems. Experimental results on the FFHQ and ImageNet datasets demonstrate state-of-the-art posterior-sampling performance, surpassing both training-based and zero-shot alternatives.  [^1] Code at https://github.com/noamelata/InvFussion.
Equal Contribution.
@@ 1 Introduction
Diffusion models [45, 48, 26] have emerged as a leading class of generative machine learning techniques [14, 40]. Since their inception, diffusion models have gained significant traction in solving complex inverse problems, such as super-resolution [43] or in-painting [35], where the goal is to reconstruct or estimate an underlying image from partial or degraded observations.
@@
Diffusion models can be trained to solve inverse problems using a simple conditioning framework, in which the measurements are supplied to the network [42, 43]. Although powerful, these methods are typically restricted to handling a limited set of degradations per trained network. This is because during training, the network must infer the connection between clean images and their degraded versions, even when a closed-form relation for each possible degradation is known. And at test time, the network must infer from the input which among the possible degradation processes generated it (a blind restoration setting). This fundamental limitation significantly reduces the applicability of trained conditional diffusion-based methods, as in many cases it is required to accommodate a wide range of inverse problems. Conversely, zero-shot methods [8, 9, 30, 31, 47] utilize unconditional pre-trained diffusion models for solving inverse problems, taking into account the particular degradation process from which each input suffers (a nonblind setting). They are thus remarkably flexible in handling various degradation scenarios. However, these approximate approaches suffer from reduced accuracy and computational inefficiency.
@@
Figure 1: Examples for posterior samples from our degradation-aware diffusion model at $256\times 256$ resolution. A single model can restore multiple degradations, such as in-painting, de-blurring and super-resolution with high image fidelity and realism, by integrating the degradation operator into the model’s architecture.
@@
Here, we introduce an architectural framework that combines the strengths of the training-based and zero-shot methodologies. Our approach relies on a new network architecture that is conditioned on the degradation operator. This conditioning enables the network to make full use of the provided measurements. Since the set of all possible degradations is too large to be used directly as a model input, we propose integrating the degradation operator into the architecture’s design, implicitly enabling the network to infer whether its intermediate features are consistent with the measurements. This novel architecture allows the network to adapt to virtually any degradation, while simultaneously maintaining the high-performance of training-based networks. Our method represents a paradigm shift, offering a more versatile, accurate, and computationally efficient approach to handling complex inverse problems.
@@
Our network’s architecture is inspired by unrolling frameworks [50, 57, 1], in which the layers mimic an optimization process. In our case, we integrate the degradation operator within the network’s layers using attention mechanisms [53, 15, 39]. The attention layers preserve dimensionality, and thus enable applying the degradation operator to all network activations. We use HDiT [11] as our base architecture, which is particularly effective in facilitating operations on higher-resolution images, as shown in Fig. 1.
Accordingly, while traditional training-based methods cannot handle scenarios where it is impossible to infer the degradation process from the measurements, and zero-shot methods are inaccurate, our model offers significant advantages in both computational efficiency and flexibility. In particular, our method can be used with any sampler for unconditional diffusion models [33, 34, 28], including Classifier-Free Guidance (CFG) [25, 10] and related techniques. Beyond posterior sampling, we illustrate how our architecture can be used for outputting the Minimum Mean Square Error (MMSE) predictor, as well as for training a Neural Posterior Principal Component (NPPC) [38] predictor, enabling uncertainly quantification for a wide variety of restoration tasks with one model.
Our experimental results on FFHQ [27] and ImageNet [13] demonstrate superior performance compared to existing inverse problem solvers, both training-based and zero-shot, establishing a new state-of-the-art (SOTA).
@@ 2.1 Diffusion Models
[…]
@@ 2.2 Inverse Problems
A popular approach for generating a reconstruction $\hat{{\mathbf{x}}}$ of ${\mathbf{x}}$ based on the measurement ${\mathbf{y}}$ and on knowledge of ${\bm{H}}$, is to draw $\hat{{\mathbf{x}}}$ from the posterior distribution,
$\hat{{\mathbf{x}}}_{\text{Post}}\sim p({\mathbf{x}}|{\mathbf{y}},{\bm{H}}).$ (2)
Methods aiming for this solution are called posterior samplers and are the main focus of this work. Another popular approach to generating a reconstruction, is by aiming for the Minimum Mean Square Error (MMSE) predictor,
$\hat{{\mathbf{x}}}_{\text{MMSE}}=\mathbb{E}\left[{\mathbf{x}}|{\mathbf{y}},{\bm{H}}\right],$ (3)
which, as its name implies, achieves the lowest possible squared-error distortion. Obtaining an approximation of the MMSE predictor can be done by training a regression network $m_{\theta}({\mathbf{y}})$ to minimize $\mathcal{L}=\mathbb{E}[\|m_{\theta}({\mathbf{y}})-{\mathbf{x}}\|^{2}]$.
@@
[…]
This posterior is a weighted average of all possible posterior distributions (for all possible degradations ${\bm{H}}$), each weighted by the corresponding $p({\bm{H}}|{\mathbf{y}})$. This posterior may generally encompass a significantly larger uncertainty regarding ${\mathbf{x}}$. One key limitation of training-based posterior samplers is that they lack a mechanism to condition the model on ${\bm{H}}$. Therefore, even if ${\bm{H}}$ is known, it cannot be provided to the model at test time, so that the model is unavoidably tasked with solving a blind restoration problem. Here, we introduce a degradation-aware model architecture, which allows solving the easier nonblind restoration task whenever the degradation is known at test time.
@@ 2.3 Diffusion Restoration
In recent years, diffusion models have become the leading approach for posterior sampling. By modifying the probability distribution into a conditional one, $p({\mathbf{x}}_{t}|{\mathbf{y}})$, solutions to the inverse problem can be sampled using the score function of these conditional distributions. The most straightforward approach to leverage diffusion models for posterior sampling is to alter the training to accommodate the partial measurements ${\mathbf{y}}$ [42, 43], modifying Eq. 1 to
@@
An alternative approach is to use an existing unconditional model $m_{\theta}({\mathbf{x}}_{t},t)$ and knowledge of ${\bm{H}}$ to model the conditional score $\nabla_{{\mathbf{x}}_{t}}\log(p({\mathbf{x}}_{t}|{\mathbf{y}},{\bm{H}}))$, using the Bayse rule
$\nabla_{{\mathbf{x}}_{t}}\log(p({\mathbf{x}}_{t}|{\mathbf{y}},{\bm{H}}))=\\ \nabla_{{\mathbf{x}}_{t}}\log(p({\mathbf{x}}_{t}))+\nabla_{{\mathbf{x}}_{t}}\log(p({\mathbf{y}}|{\mathbf{x}}_{t},{\bm{H}})).$ (6)
Such methods are referred to as zero-shot [9, 8, 31, 30, 47, 55, 56], for their use of pre-trained diffusion models to solve inverse problem tasks. These methods are highly flexible, for a single trained model can be used to solve any inverse problem for which $\nabla_{{\mathbf{x}}_{t}}\log(p({\mathbf{y}}|{\mathbf{x}}_{t},{\bm{H}}))$ can be approximated. Yet, these methods are often inaccurate, slow, and computationally expensive, due to the challenge of obtaining an accurate approximation for $\nabla_{{\mathbf{x}}_{t}}\log(p({\mathbf{y}}|{\mathbf{x}}_{t},{\bm{H}}))$. In contrast to the correctness guarantees of training-based models [4], zero-shot models cannot sample from the posterior even with an ideal denoiser [20].
@@ 2.4 Algorithm Unrolling
Algorithm unrolling, also known as deep unfolding or algorithm unrolling, represents a paradigm where iterative algorithms inspired by classical optimization utilize repeated application of neural networks. This approach, first introduced in Gregor and LeCun [19], had emerged into image processing [50, 57, 1]. In inverse problem solving, many algorithm unrolling methods utilize repeated application of ${\bm{H}}$ and ${\bm{H}}^{\dagger}$ (or ${\bm{H}}^{T}$) between network evaluations. Our InvFussion architecture is inspired by the same principles, but leads to a very different execution. Instead of approximating an iterative algorithm, it treats each of the model’s many layer’s as being unrolled through the degradation. Using the attention mechanism, the model learns where to utilize this additional information.
@@ 2.5 Attention for Conditional Generation
Attention mechanisms [53], originally developed for language processing tasks, have emerged as a dominant architectural paradigm in image processing applications. The fundamental operation in attention layers involves matching queries and keys, enabling the network to identify and combine corresponding features across different data streams, and output relevant values. This architectural approach has found widespread adoption in image generation tasks, particularly in conditional generation scenarios. A prominent example is text-to-image synthesis [41, 43], where cross-attention mechanisms facilitate interaction between textual inputs and internal diffusion model activations. The versatility of attention mechanisms extends beyond text-to-image applications, demonstrating significant utility in various domains, such as image editing [23, 52, 7], novel-view-synthesis [51, 16, 44] and many more.
@@ 3 Method
The integration of degradation operators directly as input to neural networks presents significant computational and architectural challenges. The complexity arises from the vast dimensionality of potential degradation representations. Even when considering only linear degradations, each row of the degradation matrix can be as large as the input image itself, creating an overwhelming input space that would render traditional network architectures computationally intractable and practically infeasible.
To address this fundamental challenge, our approach incorporates the degradation operator within the network architecture itself. We introduce a novel Feature Degradation Layer, which applies the degradation operator to the internal network activations, and compares the result to the provided measurements, as shown in Fig. 2. Specifically, in this layer, feature representations undergo the following operations: deep features are first rearranged in the shape of a stack or images, to which the degradation is applied. These degraded features are then concatenated with measurements, and are transformed with a learned linear operator, incorporating the information from both the measurements and the degraded features. Finally, the transformed degraded features are mapped back to the feature space through a pseudo-inverse operation. This process can be described formally as
@@
Figure 2: A diagram of the InvFussion Block. Our block contains a Feature Degradation Layer, which incorporates the operator ${\bm{H}}$ into the architecture by applying the operator on the activations and comparing them against the measurements ${\mathbf{y}}={\bm{H}}{\mathbf{x}}$.
We interweave the output of our Feature Degradation Layers back into the network’s processing stream using joint-attention, as illustrated in Fig. 2. These complete blocks, which we name InvFussion Blocks, are designed to replace standard attention blocks in existing Transformer or UNet architectures. The above operations require that the network’s deep feature resolution match the input image dimension $D$. This constraint is easy to satisfy in Vision Transformer [15, 39, 11] networks, by simply un-patching the activations back to the original image shape, and using a number of hidden channels that is divisible by the original image channels.
Finally, using our complete InvFussion architecture, a diffusion model can be trained directly by conditioning on both the degradation and the measurements,
@@
Sampling from InvFussion maintains fundamental compatibility with standard diffusion model sampling procedures, offering a significant advantage in terms of flexibility and implementation. This compatibility means that any established sampling technique developed for conventional diffusion models can be directly applied to InvFussion.

### v2

# InvFusion: Bridging Supervised and Zero-shot Diffusion for Inverse Problems

@@ Abstract
Diffusion Models have demonstrated remarkable capabilities in handling inverse problems, offering high-quality posterior-sampling-based solutions. Despite significant advances, a fundamental trade-off persists regarding the way the conditioned synthesis is employed: Zero-shot approaches can accommodate any linear degradation but rely on approximations that reduce accuracy. In contrast, training-based methods model the posterior correctly, but cannot adapt to the degradation at test-time. Here we introduce InvFusion, the first training-based degradation-aware posterior sampler. InvFusion combines the best of both worlds – the strong performance of supervised approaches and the flexibility of zero-shot methods. This is achieved through a novel architectural design that seamlessly integrates the degradation operator directly into the diffusion denoiser. We compare InvFusion against existing general-purpose posterior samplers, both degradation-aware zero-shot techniques and blind training-based methods. Experiments on the FFHQ and ImageNet datasets demonstrate state-of-the-art performance. Beyond posterior sampling, we further demonstrate the applicability of our architecture, operating as a general Minimum Mean Square Error predictor, and as a Neural Posterior Principal Component estimator. [^1] Code implementation available at https://github.com/noamelata/InvFusion
@@ 1 Introduction
Diffusion models [sohl2015deep, song2019generative, ho2020denoising] have emerged as a leading class of generative machine learning techniques [dhariwal2021diffusion, latent_diffusion]. Since their inception, diffusion models have gained significant traction in solving complex inverse problems, such as super-resolution [saharia2022image] or in-painting [lugmayr2022repaint], where the goal is to reconstruct or estimate an underlying image from partial or degraded observations.
@@
Diffusion models can be trained to solve inverse problems using a simple conditioning framework, in which the measurements are supplied to the network [saharia2022palette, saharia2022image]. Although powerful, these training-based methods are typically restricted to handling a limited set of degradations per trained network, as they must learn the connection between clean images and their degraded versions. Moreover, existing training-based models cannot accept the degradation as an input and must rely on the measurements provided at test-time to infer which among the possible degradations to restore. This forces a blind restoration setting, which is inaccurate when multiple degradations fit the measurements. These fundamental limitations significantly reduce the applicability of trained conditional diffusion-based methods, as no single model can accommodate a wide range of inverse problems. Conversely, zero-shot methods [chung2022improving, chung2023diffusion, kawar2021snips, kawar2022denoising, song2023pseudoinverseguided] utilize unconditional pre-trained diffusion models for solving inverse problems, taking into account the particular degradation process from which each input suffers, enabling a more precise non-blind setting. They are thus remarkably flexible in handling various degradation scenarios. However, these approaches use approximation that reduce accuracy and suffer from computational inefficiency.
@@
Figure 1: Examples for posterior samples from our degradation-aware diffusion model. A single model can restore multiple degradations, such as in-painting, de-blurring and super-resolution with high image fidelity and realism, by integrating the degradation operator into the model’s architecture.
@@
Here, we introduce InvFusion, an architectural framework that combines the strengths of training-based and zero-shot methodologies. Our approach is the first training-based method that can be informed of the degradation at test-time. The InvFusion architecture is conditioned on the degradation operator, enabling the network to make full use of the provided measurements. Since the set of all possible degradations is too large to be used directly as a model input, we propose integrating the degradation operator into the architecture’s design, implicitly enabling the network to infer whether its intermediate features correspond to the measurements. This novel architecture allows the network to adapt to a wide range of degradations while simultaneously maintaining the high performance of training-based networks. Our method represents a paradigm shift, offering a more versatile, accurate, and computationally efficient approach to handling complex inverse problems.
@@
Our network’s architecture is inspired by unrolling frameworks [sun2016deep, zhang2018ista, aggarwal2018modl], in which the layers mimic an optimization process. In our case, we integrate the degradation operator within the network’s layers using attention mechanisms [attention, dosovitskiy2021vit, Peebles2022DiT] . The attention layers preserve dimensionality, and thus enable applying the degradation operator to all network activations. We use HDiT [crowson2024scalable] as our base architecture, which is particularly effective in facilitating operations on higher-resolution images, as shown in Fig.˜1.
Accordingly, our method bypasses the approximations of zero-shot methods and solves the degradation-ambiguity of existing training-based method, offering significant advantages in accuracy, computational efficiency, and flexibility. In particular, our method can be used with any sampler designed for unconditional diffusion models [lu2022dpm, lu2022dpmpp, karras2022elucidating], including Classifier-Free Guidance (CFG) [ho2022classifier, chung2024cfgpp] and related techniques.
Our experiments on FFHQ [ffhq] and ImageNet [imagenet] evaluate various approaches for addressing multiple restoration tasks with a single trained model. InvFusion out-performs existing inverse problem solvers, both in the training-based and zero-shot categories, establishing a new state-of-the-art (SOTA). Beyond posterior sampling, we illustrate how our architecture can also be used for Minimum Mean Square Error (MMSE) prediction, as well as for training a Neural Posterior Principal Component (NPPC) [nehme2024uncertainty] predictor, enabling uncertainty quantification for a wide variety of restoration tasks with a single model.
@@ 2.1 Diffusion Models
Diffusion Models [sohl2015deep, song2019generative, ho2020denoising] generate high-quality images through a sequential Gaussian noise removal process. Given a source data distribution $p({\mathbf{x}}_{0})$, a forward diffusion process constructs a Stochastic Differential Equation (SDE) ${\textnormal{d}}{\mathbf{x}}_{t}=-\tfrac{1}{2}\beta(t){\mathbf{x}}_{t}+\sqrt{\beta(t)}{\textnormal{d}}{\mathbf{w}}_{t}$, where ${\mathbf{w}}_{t}$ is a standard Wiener process and $\beta(t)$ is a deterministic function of $t\in[0,1]$ taken such that at $t=1$, the final marginal distribution $p({\mathbf{x}}_{1})$ becomes approximately a standard Gaussian. Each marginal distribution along this SDE, $p({\mathbf{x}}_{t})$, can be constructed directly by adding Gaussian noise to the source data distribution, ${\mathbf{x}}_{t}=\sqrt{\bar{\alpha}(t)}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}(t)}{\mathbf{\epsilon}}$, where $\smash{\bar{\alpha}(t)=e^{-\int_{0}^{t}\beta(s)ds}}$ and $\mathbf{\epsilon}$} is a standard Gaussian random vector. To generate a sample image within this framework, one starts with a sample of white Gaussian noise and employs an appropriate SDE solver [lu2022dpm, lu2022dpmpp, karras2022elucidating] for the reverse SDE [anderson1982reverse, maoutsa2020interacting, song2020score]. The latter requires knowing the score $\nabla_{{\mathbf{x}}_{t}}\log(p({\mathbf{x}}_{t}))$ [song2019generative, song2020denoising, song2020score]. By exploiting the connection between the score and the MMSE predictor [vincent2011connection, alain2014regularized], ${\mathbf{x}}_{t}+(1-\bar{\alpha}(t))\nabla_{{\mathbf{x}}_{t}}\log(p({\mathbf{x}}_{t}))=\sqrt{\bar{\alpha}(t)}\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{t}]$, a denoising network trained to approximate $\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{t}]$ can be used in place of the real score function. Training such a network is done using a regression loss,
@@ 2.2 Inverse Problems
A popular approach for generating a reconstruction $\hat{{\mathbf{x}}}$ of $\mathbf{x}$} based on the measurement $\mathbf{y}$} and on knowledge of $\bm{H}$}, is to draw $\hat{{\mathbf{x}}}$ from the posterior distribution, $\hat{{\mathbf{x}}}_{\text{Post}}\sim p({\mathbf{x}}|{\mathbf{y}},{\bm{H}})$. Methods aiming for this solution are called posterior samplers and are the main focus of this work. Another popular approach to generating a reconstruction, is aiming for the Minimum Mean Square Error (MMSE) predictor, $\hat{{\mathbf{x}}}_{\text{MMSE}}=\mathbb{E}\left[{\mathbf{x}}|{\mathbf{y}},{\bm{H}}\right]$, which, as its name implies, achieves the lowest possible squared-error distortion. Obtaining an approximation of the MMSE predictor can be done by training a regression network $m_{\theta}({\mathbf{y}})$ to minimize $\mathcal{L}=\mathbb{E}[\|m_{\theta}({\mathbf{y}})-{\mathbf{x}}\|^{2}]$.
@@
$p({\mathbf{x}}|{\mathbf{y}})=\int p({\mathbf{x}}|{\mathbf{y}},{\bm{H}})p({\bm{H}}|{\mathbf{y}})d{\bm{H}}.$ (2)
This posterior is a weighted average of all possible posterior distributions (for all possible degradations $\bm{H}$}), each weighted by the corresponding $p({\bm{H}}|{\mathbf{y}})$. This posterior may generally encompass a significantly larger uncertainty regarding $\mathbf{x}$}. One key limitation of training-based posterior samplers (like Palette [saharia2022palette]) is that they currently lack a mechanism to condition the model on $\bm{H}$}. Therefore, even if $\bm{H}$} is known, it cannot be provided to the model at test time, so that the model is unavoidably tasked with solving a blind restoration problem. In this work, we introduce a degradation-aware model architecture, enabling the model to incorporate information about the degradation $\bm{H}$}. This allows solving the precise non-blind restoration task whenever the degradation is known at test time.
@@ 2.3 Diffusion Restoration
In recent years, diffusion models have become the leading approach for posterior sampling. By modifying the probability distribution into a conditional one, $p({\mathbf{x}}_{t}|{\mathbf{y}})$, solutions to the inverse problem can be sampled using the score function of these conditional distributions. The most straightforward approach to leverage diffusion models for posterior sampling is to alter the training to accommodate the partial measurements $\mathbf{y}$} [saharia2022palette, saharia2022image], modifying Eq.˜1 to
@@
An alternative approach is to use an existing unconditional model $m_{\theta}({\mathbf{x}}_{t},t)$ and knowledge of $\bm{H}$} to model the conditional score $\nabla_{{\mathbf{x}}_{t}}\log(p({\mathbf{x}}_{t}|{\mathbf{y}},{\bm{H}}))$, using the Bayes rule
$\nabla_{{\mathbf{x}}_{t}}\log(p({\mathbf{x}}_{t}|{\mathbf{y}},{\bm{H}}))=\nabla_{{\mathbf{x}}_{t}}\log(p({\mathbf{x}}_{t}))+\nabla_{{\mathbf{x}}_{t}}\log(p({\mathbf{y}}|{\mathbf{x}}_{t},{\bm{H}})).$ (4)
Such methods are referred to as zero-shot [chung2023diffusion, chung2022improving, kawar2022denoising, kawar2021snips, song2023pseudoinverseguided, wang2022zero, zhang2024improving], for their use of pre-trained diffusion models to solve inverse problem tasks. These methods are highly flexible, for a single trained model can be used to solve any inverse problem for which $\nabla_{{\mathbf{x}}_{t}}\log(p({\mathbf{y}}|{\mathbf{x}}_{t},{\bm{H}}))$ can be approximated. Yet, these methods are often inaccurate, slow, and computationally expensive, due to the challenge of obtaining an accurate approximation for $\nabla_{{\mathbf{x}}_{t}}\log(p({\mathbf{y}}|{\mathbf{x}}_{t},{\bm{H}}))$. In contrast to the correctness guarantees of training-based models [batzolis2021conditional], zero-shot models cannot sample from the posterior even with an ideal denoiser [gupta2024diffusion].
@@ 2.4 Algorithm Unrolling
Algorithm unrolling, also known as deep unfolding, represents a paradigm where iterative algorithms inspired by classical optimization utilize repeated application of neural networks. This approach, first introduced in gregor2010learning, had emerged into image processing [sun2016deep, zhang2018ista, aggarwal2018modl, monga2021algorithm]. In inverse problem solving, many algorithm unrolling methods apply $\bm{H}$} and $\bm{H}$$^{\dagger}$ (or ${\bm{H}}^{T}$) between network evaluations, and are trained in an end-to-end manner. Deep unrolling algorithms are typically trained for distortion reduction and do not accommodate posterior sampling, unlike our method, which is a diffusion model. Also, while the InvFusion architecture is inspired by the same principles, it uses a very different execution. Instead of approximating an iterative algorithm, InvFusion uses the attention mechanism to learn where to utilize the knowledge of the degradation.
@@ 2.5 Attention for Conditional Generation
Attention mechanisms [attention], originally developed for language processing tasks, have emerged as a dominant architectural paradigm in image processing applications. The fundamental operation in attention layers involves matching queries and keys, enabling the network to identify and combine corresponding features across different data streams, and output relevant values. This architectural approach has found widespread adoption in image generation tasks, particularly in conditional generation scenarios. A prominent example is text-to-image synthesis [rombach2022high, saharia2022image], where cross-attention mechanisms facilitate interaction between textual inputs and internal diffusion model activations. The versatility of attention mechanisms extends beyond text-to-image applications, demonstrating significant utility in various domains, such as image editing [hertz2022prompt, Tumanyan_2023_PNP, cao_2023_masactrl], novel-view-synthesis [tseng2023consistent, elata2024novel, seo2024genwarp] and many more.
@@ 3 Method
The integration of degradation operators directly as input to neural networks presents significant computational and architectural challenges. The complexity arises from the vast dimensionality of potential degradation representations. Even when considering only linear degradations, each row of the degradation matrix is as large as the input image itself, creating an overwhelming input space that would render traditional network architectures computationally intractable and practically infeasible.
To address this fundamental challenge, our approach incorporates the degradation operator within the network architecture itself. We introduce a novel Feature Degradation Layer, which applies the degradation operator to the internal network activations, and compares the result to the provided measurements, as shown in Fig.˜2.
Specifically, in this layer, feature representations undergo the following operations: deep features are first rearranged in the shape of a stack of images, to which the degradation is applied. These degraded features are then concatenated with the measurements, and are transformed with a learned linear operator, incorporating the information from both the measurements and the degraded features.
Figure 2: A diagram of the InvFusion Block. Our block contains a Feature Degradation Layer, which incorporates the operator $\bm{H}$} into the architecture by applying $\bm{H}$} on the activations and comparing them with the measurements ${\mathbf{y}}={\bm{H}}{\mathbf{x}}$.
Finally, the transformed degraded features are mapped back to the feature space through the degradation’s pseudo-inverse. This process can be described formally as
@@
We interweave the output of our Feature Degradation Layers back into the network’s processing stream using joint-attention [^3] Joint-attention differs from cross-attention by using the same queries for both the “self” and “cross” keys, as illustrated in Fig.˜2. These complete blocks, which we name InvFusion Blocks, are designed to replace some or all standard attention blocks in existing Transformer or UNet architectures. The above operations require that the network’s deep feature resolution match the input image dimension $D$. This constraint is readily accommodated in many Vision Transformer [dosovitskiy2021vit, Peebles2022DiT, crowson2024scalable] based networks by using a number of hidden channels that is divisible by the original image channels (which does not compromise performance). Internal network activations can then be un-patched back to the original image shapes to apply the degradation $\bm{H}$}.
Finally, using our complete InvFusion architecture, a diffusion model can be trained directly by conditioning on both the degradation and the measurements,
@@
Sampling from InvFusion maintains fundamental compatibility with standard diffusion model sampling procedures, offering a significant advantage in terms of flexibility and implementation. This compatibility means that any established sampling technique developed for conventional diffusion models can be directly applied to InvFusion.
3.1 MMSE Estimator
Our InvFusion architecture can be used for more than posterior sampling. Having an architecture that makes use of the degradation $\bm{H}$} enables training a single model with
to produce an accurate MMSE predictor for many different degradations. Alternatively, the MMSE predictor can be approximated by a diffusion model trained with Eq.˜6 by applying a single denoising step on white Gaussian noise. That is, for sufficiently small $\bar{\alpha}(t)$ we have that ${\mathbf{x}}_{t}\approx{\mathbf{\epsilon}}$, which is independent of ${\mathbf{x}}_{0}$, so that
which also holds for the approximate models being trained. In other words, any conditional diffusion model can be used as an efficient MMSE predictor. We find that this works well (see App.˜B), and utilize this in our experiments.